OpenStudy (themathasker):

HELP WILL FAN AND MEDAL! On a right triangle, how do I use the Pythagorean theorem to find b=10 c=25 and a=?

OpenStudy (4everaddicted2anime):

the formula is a^2 + b^2 = c^2 so you need to plug into the values for each variable. a^2 + 10^2 = 25^2

OpenStudy (4everaddicted2anime):

You need to simplify the problem by squaring what numbers can be squared. a^2 + 100 = 625 Next you need to isolate the a variable by subtracting 100 from both sides. a^2 = 525 Now you need to take the square root of both sides to get your answer for a. a=22.91

OpenStudy (themathasker):

I understand now! Thanks so much!!!!!!! :D

OpenStudy (4everaddicted2anime):

your welcome